OpenStudy (animelover91804):

1. The ratio of teachers to students is 1 to 26. Find all equivalent ratios. * a. 52 to 2 b. 11 to 286 c. 26 to 1 d. 5 to 130 e. 0.25 to 6.5 f. 1/2 to 19 1/2 2. Express the ratio in simplest form by first converting to same unit of measure : 3 inches to 2 feet. * 3 Express the rate as a unit rate: $258 for 6 tires. *

OpenStudy (anonymous):

B,D, E (:

OpenStudy (anonymous):

B. because 11*26=286 and 1*11=11 D. because 5*26=130 and 1*5=5 E. because 26*.25=6.5 and 1*.25=.25

OpenStudy (anonymous):

2. You would have to get the feet into inches. Since there is 12 inches in a foot, 12 x 2 = 24. So, 3 to 24. Simplify by dividing by 3, 1 to 8. So the answer is: 1 inch to 8 inches

OpenStudy (animelover91804):

5.A 2006 Harley-Davidson Fat-Boy Motorcycle can travel 69 miles on 1 gallons of gasoline. Write and solve a proportion to determine how many gallons of gasoline this motorcycle will need to travel 350 miles. (Round answer to nearest tenth.)